A capacitor with an initial potential difference of 117 V is discharged through a resistor when a switch between them is closed at t = 0 s. At t = 10.0 s, the potential difference across the capacitor is 1.11 V. (a) What is the time constant of the circuit? (b) What is the potential difference across the capacitor at t = 15.7 s?
Answer #1

Equation for potential across the capacitor while discharging is

VC = Vo e-t/RC   where RC is the time constant.

Vo is the initial voltage which is given as 117 V

VC at t = 10 sec is 1.11 V

So, 1.11 = 117 e-10/RC

1.11/117 =  e-10/RC

-4.657813919 = -10/RC

RC = 2.146929906 seconds (answer for part a)

VC at t = 15.7 seconds :

VC = 117 e-15.7/2.146929906 = 0.078035305

therefore VC at t= 15.7 seconds is 0.078 Volts